Advertisements
Advertisements
प्रश्न
A problem in Mathematics is given to three students whose chances of solving it are `1/3, 1/4` and `1/5`. What is the probability that exactly one of them will solve it?
Advertisements
उत्तर
`"P"(bar"A"_1) = 1 - "P"("A"_1) = 1 - 1/3 = 2/3`
`"P"(bar"A"_1) = 1 - "P"("A"_2) = 1 - 1/4 = 3/4`
`"P"(bar"A"_1) = 1 - "P"("A"_3) = 1 - 1/5 = 4/5`
Probability of Exactly one student solving the problem
= `"P" [("A"_1 ∩ bar"A"_2 ∩ bar"A"_3) ∪ (bar"A"_1 ∩ "A"_2 ∩ bar"A"_3) ∪ (bar"A"_1 ∩ bar"A"_2 ∩ "A"_3)]`
= `"P"("A"_1 ∩ bar"A"_2 ∩ bar"A"_3) + "P"(bar"A"_1 ∩ "A"_2 ∩ bar"A"_3) + "P"(bar"A"_1 ∩ bar"A"_2 ∩ "A"_3)`
= `"P"("A"_1) "P"(bar"A"_2) "P"(bar"A"_3) + "P"(bar"A"_1) "P"("A"_2) "P"(bar"A"_3) + "P"(bar"A"_1) "P"(bar"A"_2) "P"("A"_3)`
= `1/3 xx 3/4 xx 4/5 + 2/3 xx 1/4 xx 4/5 + 2/3 xx 3/ xx 1/5`
= `(12 + 8 + 6)/60`
= `26/16`
= `13/30`
[Probability of exactly one student solving the problem = Probability [(A1 solving the problem and A2, A3 non solving the problem) or (A1, A3 non solving and A2 solving) or (A1, A2 solving and A3 non solving)]
APPEARS IN
संबंधित प्रश्न
If P(A) = 0.8, P(B) = 0.5 and P(B|A) = 0.4, find
- P(A ∩ B)
- P(A|B)
- P(A ∪ B)
If P(A) = 0.8, P(B) = 0.5 and P(B|A) = 0.4, find P(A|B)
Determine P(E|F).
A coin is tossed three times, where
E: at most two tails, F: at least one tail
Determine P(E|F).
Two coins are tossed once, where
E: tail appears on one coin, F: one coin shows head
Determine P(E|F).
Two coins are tossed once, where
E: no tail appears, F: no head appears
A fair die is rolled. Consider events E = {1, 3, 5}, F = {2, 3} and G = {2, 3, 4, 5} Find P (E|F) and P (F|E)
A fair die is rolled. Consider events E = {1, 3, 5}, F = {2, 3} and G = {2, 3, 4, 5} Find P ((E ∪ F)|G) and P ((E ∩ G)|G)
An instructor has a question bank consisting of 300 easy True/False questions, 200 difficult True/False questions, 500 easy multiple choice questions and 400 difficult multiple choice questions. If a question is selected at random from the question bank, what is the probability that it will be an easy question given that it is a multiple-choice question?
Suppose we have four boxes. A, B, C and D containing coloured marbles as given below:
| Box | Marble colour | ||
| Red | White | Black | |
| A | 1 | 6 | 3 |
| B | 6 | 2 | 2 |
| C | 8 | 1 | 1 |
| D | 0 | 6 | 4 |
One of the boxes has been selected at random and a single marble is drawn from it. If the marble is red, what is the probability that it was drawn from box A?, box B?, box C?
One bag contains 5 white and 3 black balls. Another bag contains 4 white and 6 black balls. If one ball is drawn from each bag, find the probability that both are black
Choose the correct alternative:
A letter is taken at random from the letters of the word ‘ASSISTANT’ and another letter is taken at random from the letters of the word ‘STATISTICS’. The probability that the selected letters are the same is
If X denotes the number of ones in five consecutive throws of a dice, then P(X = 4) is ______
Let A and B be two events. If P(A) = 0.2, P(B) = 0.4, P(A ∪ B) = 0.6, then P(A|B) is equal to ______.
If P(A) = 0.4, P(B) = 0.8 and P(B|A) = 0.6, then P(A ∪ B) is equal to ______.
Let A and B be two non-null events such that A ⊂ B. Then, which of the following statements is always correct?
If A and B are two independent events such that P(A) = `1/3` and P(B) = `1/4`, then `P(B^'/A)` is ______.
If A and B are any two events of S and \[P(F)\neq 0\], which formula is correct?
Which statement correctly describes the numerator in \[P(A\mid B)=\frac{P(A\cap B)}{P(B)}\]?
